OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5). −1 1 0

OpenStudy (anonymous):

-1

OpenStudy (hoblos):

slope = (-5+3)/(3-1) = -2/2 =-1

OpenStudy (anonymous):

\[-5 + 3 \over 3 - 1\] =